$SPAD/src/input drawcx.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

\getchunk{license}
)clear all
--Plotting Two Dimensional Compiled Functions of One Variable
f(x:DFLOAT):DFLOAT == sin(tan(x))-tan(sin(x))
draw(f,0..6)
draw(f,0..6,unit == [1.0,0.5],toScale == true)
--Plotting Two Dimensional Compiled Parametric Plane Curve Functions
i1(t:DFLOAT):DFLOAT == 9*sin(3*t/4)
i2(t:DFLOAT):DFLOAT == 8*sin(t)
draw(curve(i1,i2),-4*%pi..4*%pi)
draw(curve(i1,i2),-4*%pi..4*%pi,toScale == true, title == "Lissajous Curve")
--Plotting Three Dimensional Compiled Parametric Space Curve Functions
i1(t:DFLOAT):DFLOAT ==sin(t)*cos(3*t/5)
i2(t:DFLOAT):DFLOAT ==cos(t)*cos(3*t/5)
i3(t:DFLOAT):DFLOAT ==cos(t)*sin(3*t/5)
draw(curve(i1,i2,i3),0..15*%pi)
draw(curve(i1,i2,i3),0..15*%pi,title == "Parametric Curve")
--Plotting Three Dimensional Compiled Functions of Two Variables
l(x:DFLOAT,y:DFLOAT):DFLOAT == cos(x*y)
draw(l,-3..3,-3..3)
colorFxn(x:DFLOAT,y:DFLOAT):DFLOAT == 1/(x**2 + y**2 + 1)
draw(l,-3..3,-3..3, colorFunction == colorFxn)
--Plotting Three Dimensional Compiled Parametric Surface Functions
n1(u:DFLOAT,v:DFLOAT):DFLOAT == u*cos(v)
n2(u:DFLOAT,v:DFLOAT):DFLOAT == u*sin(v)
n3(u:DFLOAT,v:DFLOAT):DFLOAT == v*cos(u)
draw(surface(n1,n2,n3),-4..4,0..2*%pi)
colorFxn(x:DFLOAT,y:DFLOAT):DFLOAT == 1/(x**2 + y**2 + 1)
draw(surface(n1,n2,n3),-4..4,0..2*%pi, colorFunction == colorFxn)

———-

References^{[1] nothing}

4